OpenStudy (anonymous):

How does expression written in exponential form: 6^log6(x) turn to a logarithm form: 6^log6(x)=x? HELP!

OpenStudy (anonymous):

this is property of log \[\Huge a^{\log_{a}(x)}=x\] where a here is a=6

OpenStudy (anonymous):

is it something like: 6^x = y 6^y= x... and then I'm stuck

OpenStudy (anonymous):

you can write 6^x=y as \[\Large \log_{6} y=x\]

OpenStudy (anonymous):

on my textbook it states, 6^log6(x) 6^log6(x) = y log6y = log6x y=x 6^log6(x) = x

OpenStudy (anonymous):

but in the question the log was in the exponent of the base of log 6 so they usualyy gets canceled

OpenStudy (anonymous):

i don't understand why... :(

OpenStudy (anonymous):

ohh.. can you please show me a step-by-step? :$